OpenStudy (anonymous):

Let L be the line of intersection of the two planes: 3x-7y+2z=23 x-2y+3z=6 let L' be the line through the two points A=(2,-1,0) and B=(4,3,-1) (i) Find in parametric form an expression for the general point of L. (ii) Write down the vector AB and an expression for the general point of L' (iii) Determine the point in which L' meets the plane x + y + z = 6. (iv) Decide whether L meets L' (v) Decide whether L 'meets the line of intersection M of the two planes x + y + z = 6, x + 2y − 3z = −3 I cant even do the first part i dont quite get how to go about doing it. Any help is gr8

ganeshie8 (ganeshie8):

every line in a plane is perpendicular to the plane's normal vector, yes ?

OpenStudy (anonymous):

yes. I think ive done the first part

ganeshie8 (ganeshie8):

good, you just need direction vector and a point to write out the equation of line

ganeshie8 (ganeshie8):

part ii should be pretty straightforward ?

OpenStudy (anonymous):

yeah i found the vector AB just by doing a-b...but then im not sure what it wanted by the general point of L'?

ganeshie8 (ganeshie8):

vector AB is B-A

OpenStudy (anonymous):

sorry yes thats what i meant ><

ganeshie8 (ganeshie8):

just write out the equation of line

OpenStudy (jhannybean):

First you should set your functions as \(\sf z(x,y)= ...\)

OpenStudy (jhannybean):

Then for (i) your hint is: \(\sf \vec r(t) = (1-t)\vec r_0 +t\cdot \vec r_1\)

OpenStudy (jhannybean):

create parametric equations first, always always always.

OpenStudy (jhannybean):

So if you FIRST solve your equations in terms of z, then you can set them equal to each other to find the intersection :) That's what I think.

OpenStudy (anonymous):

i did that, i have my equations in terms of z

OpenStudy (jhannybean):

this looks like what I just learned about line integrals and Green's Theorem...all that yummy stuff.

OpenStudy (anonymous):

its ii), i found AB but im not sure how to find the general point of L'

ganeshie8 (ganeshie8):

are you done with part i ?

ganeshie8 (ganeshie8):

whats the equation of L ?

OpenStudy (anonymous):

i believe so, i got an answer at least

OpenStudy (anonymous):

Well i solved it in terms of Z, so i have x=7z+11 and y=-7-5z

ganeshie8 (ganeshie8):

that wont work

ganeshie8 (ganeshie8):

you need equation of form : L = `point` + t( `direction vector`)

OpenStudy (anonymous):

i reduced using REF to eliminate Z

OpenStudy (jhannybean):

And so when you set them equal to eachother you'll find the intersection? this is my guess <.<

OpenStudy (jhannybean):

What is REF...

OpenStudy (anonymous):

Row echelon form

ganeshie8 (ganeshie8):

3x-7y+2z=23 x-2y+3z=6 cross product of (3, -7, 2) and (1, -2, 3) gives you the direction vector of line

OpenStudy (jhannybean):

because (3,-7,2) and (1, -2, 3) are the normal vectors of both the planes.

OpenStudy (jhannybean):

Taking their cross product gives you a direction vector.... :D

ganeshie8 (ganeshie8):

yes start by finding the cross product jbn

OpenStudy (anonymous):

(-25,-7,1)

ganeshie8 (ganeshie8):

try again

OpenStudy (anonymous):

(-17,-7,1)

OpenStudy (jhannybean):



OpenStudy (jhannybean):

\[\sf<-21 +4 ~, -(9-2)~, -6+7>\]

ganeshie8 (ganeshie8):

right, so we have direction vector of the line : L = `point` + t `<-17, -7, 1>`

ganeshie8 (ganeshie8):

just find a `point` on the line and you're done with part i

ganeshie8 (ganeshie8):

any ideas on how to find a point on the line ?

OpenStudy (anonymous):

set them equal and solve for where they intersect?

ganeshie8 (ganeshie8):

that would be tedious

ganeshie8 (ganeshie8):

you just need some point

ganeshie8 (ganeshie8):

so a better idea is to plugin z = 0 and solve the system : 3x-7y+2(0)=23 x-2y+3(0)=6

ganeshie8 (ganeshie8):

http://www.wolframalpha.com/input/?i=solve+3x-7y%3D23%2Cx-2y%3D6 so a point on the line is `(-4, -5, 0)`

OpenStudy (jhannybean):

Or rather, set two variables = 0 and solve 1 at a time, i.e x=y=0, solve for z, etc.

ganeshie8 (ganeshie8):

parametric form of line would be : L = `<-4, -5, 0>` + t `<-17, -7, 1>`

OpenStudy (anonymous):

That was very simple, thank you

OpenStudy (anonymous):

Similar method for ii i presume?

ganeshie8 (ganeshie8):

yes u just need a `point` and `direction vector` for writing out parametric form of line

ganeshie8 (ganeshie8):

find them

ganeshie8 (ganeshie8):

part ii `let L' be the line through the two points A=(2,-1,0) and B=(4,3,-1)` direction vector = AB = `<2, 4, -1>` yes ?

ganeshie8 (ganeshie8):

you have two points already, just pick one and write out the equation

OpenStudy (anonymous):

yes thats what i got

OpenStudy (anonymous):

use a different parameter to L like s yes?

ganeshie8 (ganeshie8):

you can use ur fav letter for parameter, but using t is also okay if you're aware that it is not the same as the t in L

OpenStudy (anonymous):

okay gotcha, onto iii i go

ganeshie8 (ganeshie8):

so whats ur equation for L' we need to use it in part iii

OpenStudy (anonymous):

(2,-1,0)+s(2,4,-1)

ganeshie8 (ganeshie8):

looks good, write it in point form for part iii

ganeshie8 (ganeshie8):

L' = (2,-1,0)+s(2,4,-1) which is same as L' = (2+2s, -1+4s, -s) yes ?

OpenStudy (anonymous):

yes, and then sub the x y and z values into the equation of the plane?

ganeshie8 (ganeshie8):

plug this point in the given plane ` x + y + z = 6`

ganeshie8 (ganeshie8):

yes sub in and find the value of s at which the line and plane meet

OpenStudy (anonymous):

i get s to be 2

OpenStudy (anonymous):

so then sub this back into L' to find the point

ganeshie8 (ganeshie8):

try again

ganeshie8 (ganeshie8):

im getting s=1

OpenStudy (anonymous):

yes s=6

OpenStudy (anonymous):

sorry, 1

OpenStudy (anonymous):

so now sub the correct s( 1 not 2) back into the L'

ganeshie8 (ganeshie8):

yes

OpenStudy (anonymous):

L'=(14,23,-5)

ganeshie8 (ganeshie8):

L' = (2+2s, -1+4s, -s) sub in s=1 for the intersection point of line L' and the plane `x+y+z=6`

OpenStudy (anonymous):

i keep putting wrong values in ><

OpenStudy (anonymous):

L'=(4,3,-5)

ganeshie8 (ganeshie8):

try again

OpenStudy (anonymous):

(4,3,-1)

ganeshie8 (ganeshie8):

looks good

OpenStudy (anonymous):

so now to see if L' and L meet, i would set them equal and solve for each parameter, and if they do meet when i sub them back in i would get the same answer. yes?

ganeshie8 (ganeshie8):

yes L = (-4-17t, -5-7t, t) L' = (2+2s, -1+4s, -s) set the components equal to each other, you will get 3 equations with 2 unknowns. if the solution exists, the lines meet. otherwise they are skew

OpenStudy (anonymous):

Im getting that they dont meet

ganeshie8 (ganeshie8):

thats right, the lines done meet and are not parallel. so they are skew http://www.wolframalpha.com/input/?i=solve+-4-17t%3D2%2B2s%2C-5-7t%3D-1%2B4s%2Ct%3D-s

OpenStudy (anonymous):

So for the last part, i need to make another parametric for the intersection of M

ganeshie8 (ganeshie8):

yes find the parametric form of M first

ganeshie8 (ganeshie8):

then change both M and L' to point forms and set their components equal

ganeshie8 (ganeshie8):

solve them.. same story